Suppose your firm is considering two mutually exclusive, required projects with the cash flows shown below. The required rate of return on projects of both of their risk class is 10 percent, and that the maximum allowable payback and discounted payback statistic for the projects are 2 and 3 years, respectively.

  Time: 0 1 2 3
  Project A Cash Flow -37,000 27,000 47,000 18,000
  Project B Cash Flow -47,000 27,000 37,000 67,000


Use the NPV decision rule to evaluate these projects; which one(s) should it be accepted or rejected?

Multiple Choice

  • accept neither A nor B

  • accept A, reject B

  • accept both A and B

  • reject A, accept B
Answer #1

reject A, accept B

NPV = -Initial investment + Present value of cash flows

Project A:

NPV = -37,000 + 27,000/1.10 + 47,000/1.10^2 + 18,000/1.10^3 = 39,912.10

Project B:

NPV = -47,000 + 27,000/1.10 + 37,000/1.10^2 + 67,000/1.10^3 = 58,462.06